OpenStudy (anonymous):

Show the formula 1. The perimeter of a rectangle is 52. Its length is2 more than twice the measure of its width. Find its dimensions. 2. If the diameter of a circle is the same as the radius of the circumference of the larger circle to the circumference of the smaller.

OpenStudy (anonymous):

Alright, do you know the formula for a perimeter?

OpenStudy (anonymous):

@killky28 Are you here?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay. Do you know the formula for the perimeter of a rectangle?

OpenStudy (anonymous):

yes its l(2) + w(2)

OpenStudy (anonymous):

Yeah or just P = 2(l + w) haha okay so we have 52 = 2(l + w)

OpenStudy (anonymous):

so it will be 52= 2l + 2w

OpenStudy (anonymous):

yes :) and then hold on lol i'm writing it out

OpenStudy (anonymous):

ok ok ill wait ma'am hehe =D

OpenStudy (anonymous):

xD You don't have to call me ma'am haha I'm just a baby xD lol jk about that part

OpenStudy (anonymous):

Okay I got it!